Suppose a consumer has the following utility function:

u(x,y) = x??(1-?)

Price o f x= $3, price of ? = $7, and ? (alpha) = 0,3.

a) what is the Hicksian demand for goods x and y? Use the LaGrange method or MRS.

b) Draw the graph of the minimization (budget constraint).

Please provide step behind the solution. Thanks!
Answer #1

a)

min Sit Pxx+ Pyy U= xx y 1-2 since Px = 3 Py= 7, 2=0.3 i min 3x+74 S. 4. u= X0-340-7 L= 8L 3x+74+ lon xo 3 yo 7) 3+ - 0.3 x 00. 34-0.7 your 0.7 x x 0.3 y-03 . y = r in egu110 u= va x0.3 yo. 7 x0.3 x 6-7 . xh = u yh zu(B) Budget constraint PxX+ Pyy € 1 3x + 74 & 1 x=0, y = yao, x = 7 I stire At10